OpenStudy (anonymous):

Can anyone please explain how to use Double and Half Angle formulas to find exact values?

OpenStudy (anonymous):

of what?

OpenStudy (anonymous):

in trigonometry !

OpenStudy (anonymous):

of any given number both on and off the Unit Circle. For example if I needed to find the a) sin(2x) b) cos(2x) c) sin(x/2) and d) (cos x/2) if the cos(x)=(3/5), 0<x<(π/2)

OpenStudy (anonymous):

\[\sin (2x)=2\sin x.\cos x\] \[\cos (2x)=\cos ^{2}x -\sin ^{2}x\] in which the angle 2x is divided by 2 .. if it was x/2 << divide it bye 2 to be x/4